# Cultural Transmission, Technology, and Treatment of the Elderly

### v1

@@ 4.4 Savings, Capital Accumulation, and Inculcation
We now analyze a version of the model in which assets are accumulated via conscious savings decisions, and investigate how accumulation interacts with inculcation and gifts to the elderly. This allows us to study how inculcative investment interacts with savings and also allows us to develop the exposition in a way that is relatable to classical growth models with technological progress.
@@
The solution of (35) presents two challenges. First, a solution must take into account that current middle-aged agents’ future gifts depend on current inculcative investment through the parameter $\eta_{t+1}$, and in an equilibrium solution to the model, beliefs about this dependence must be correct and consistent with current actions. Second, a solution must also admit the possibility of corner solutions in which $\eta^{*}_{t}=0$. The subgame-perfect Nash equilibrium solution to the problem (35) has a structure in which agents give gifts in proportion to the present discounted value of lifetime income, with this proportion depending on $\eta_{t}$. For the time being, we have bypassed the issue of the determination of $\eta_{t}^{*}$ itself; as an equilibrium object, its determination can only be addressed in a fully specified general equilibrium setting. The exact form of the equilibrium gift-giving function is given in Proposition 4:
@@
See Appendix B. ∎
Equation (37) in Proposition 5 describes the total amount of current income middle-aged agents wish to allocate to future consumption. This condition shows that expenditures on savings and inculcation add up to the fraction of income a middle-aged agent wishes to save, which depends upon current income less discounted future income. Equation (38) in Proposition 5 requires that the rates of return on the different means of transferring current wealth to the future in (37) be equilibrated. That is, the rate of return on inculcative investments and what can be called conventional savings must be equal in equilibrium; otherwise, middle-aged agents would not employ both savings instruments. As we maintain the assumption that $d_{t}$ is proportional to income ($d_{t}=\delta y_{mt}$), we can interpret (38) to mean that returns to accumulated assets such as capital must be equal in equilibrium to the unit return on inculcative investment, which weighs future gifts against current inculcative costs.
The question remains whether an equilibrium with gifts and inculcation is sustainable, as both Propositions 4 and 5 do not allow much further analysis of the nature and magnitude of inculcative investments, and also proceed following the assumption that $\eta_{t}>0$. To go further, we need to round out the model to see how $\eta_{t}$, $s_{t}$, and $R_{t}$ depend on the underlying nature of the economy, as they are quantities determined in equilibrium and require specification of the production side of the economy.
@@ 4.4.1 Production and Economic Growth
To fully study the model, we need to specify the underlying nature of production in the economy. This will allow us to turn the two equations in Proposition 5 into dynamic laws describing how the effective capital stock progresses and how inculcative investments proceed. We again use a simple Cobb-Douglas production technology in which production occurs according to the aggregate production function $Y_{t}=L_{t}^{\alpha}K_{t}^{1-\alpha}$, where $K_{t}$ now denotes the total capital stock and as before $L_{t}$ denotes total effective labor supply:
@@
We assume that $A_{it},i=e,m$ grows at a constant exogenous rate $a$, while the population grows at a rate $n$. As in previous sections, constant population growth means that $N_{mt}=(1+n)N_{et}$, and that $\tfrac{1}{1+n}$ is a measure of the relative size of the elderly population.
We use (39) and the production technology to obtain the following expressions for middle-aged labor income, elderly labor income, and the gross rate of return on capital, under the maintained assumption that all resources are paid their marginal products, where $k_{t}=\frac{K_{t}}{L_{t}}$ is capital per effective unit of labor:
@@
Applying these equations to the equilibrium relationships (37) and (38) in Proposition 5 admits the description of the equilibrium paths of $\eta$ and $k$. Pursuant to analyzing (40), we also define some new terms:
@@
See Appendix C. ∎
Equation (43) describes the dynamics of inculcation but can also be viewed as a condition that describes whether or not there will be any inculcative investment at all, and therefore allows us to delineate when inculcative equilibria will be sustainable. When the right-hand side of (43) is nonpositive, $\eta_{t+1}=0$ in equilibrium. One can see that this possibility depends on the rates of population growth and technological progress; in particular, a higher rate of technological progress makes $\eta=0$ more likely. This result occurs because higher growth rates increase the demand and hence returns on capital, making inculcation less desirable. Also making $\eta=0$ more likely in equilibrium are higher inculcative costs, as measured by $\delta$, and greater direct participation in the economy by the elderly (larger values for both $\tau_{e}$ and $\gamma_{e}$).
A value of the relative consumption ratio of the elderly $\frac{c_{et}}{c_{mt}}$ can be found using the equilibrium values of savings, inculcation, and gifts. In Appendix D, we describe the calculations that lead to Proposition 7:
@@
Equation (44) in Proposition 7 indicates that the relative consumption ratio depends on relative lifetime incomes, and therefore the overall growth rate of the economy. The expression also suggests that many of the main general equilibrium implications of the model will flow through equilibrium asset returns, $R$, and the magnitude of $\eta$. This is particularly true along a balanced growth path, as in a balanced growth equilibrium, incomes grow at a rate $a$, and asset returns are constant. Along a balanced growth path, (44) reduces to:
@@
Equation (46) therefore indicates that parameters such as $\alpha$, the labor intensity of production, $n$, the growth rate of population, and $\delta$, the per-unit-income costs of inculcation will influence the relative consumption of the elderly by exerting an impact on the level of inculcation $\eta$ and equilibrium asset returns $R$. Of course, $\eta$ and $R$ are somewhat complicated functions of model parameters. We will postpone discussion of the complete balanced-growth solution of the model until Section 4.5.2.
@@ 4.5.1 Asset income as the sole source of elderly income
Equation (51) shows how the possibility of inculcation and gifts impacts the canonical overlapping-generations model with population and technological progress, as developed in, for example, Acemoglu (2009). But for the $\delta\eta_{t+1}$ term in (51), the accumulation condition is a standard capital accumulation condition in growth models, so the possibility of inculcative investment reduces the amount of capital per effective unit of labor, which will increase equilibrium asset return rates. Dividing (50) by (51) and solving the result for $\eta_{t+1}$ gives a time-independent solution for the inculcative parameter $\eta$:
@@
Equation (52) indicates that a larger $\delta$, not surprisingly, reduces the equilibrium size of $\eta^{*}$. An increase in the relative intensity of capital in production, measured by $\frac{1-\alpha}{\alpha}$, also negatively impacts $\eta$. Societies that employ more capital-intensive means of production, other things equal, will have lower degrees of inculcation and perhaps transmission of culture.
If parameters are configured so that $\eta^{*}\leq 0$, in equilibrium $\eta=0$, and there is no inculcative investment. The elderly ensure their livelihood as they would in the typical overlapping generations model: through accumulation of productive assets during middle age. Setting (52) to zero and solving for $\alpha$ gives an interpretation of (52) as a condition for inculcation to occur in terms of the relative capital intensity of production:
@@
This critical threshold depends positively on the costs of inculcation, and on $\beta$, which influences the relative weight placed on consumption in old age in utility. If (53) is not satisfied and no inculcation occurs, we recover the law of motion for the capital stock by using $\eta=0$ in (51). In forming the relative ratio of consumption, we find it most straightforward to solve the model in terms of $R$, equilibrium asset returns, rather than in terms of $k$, the equilibrium per-unit-labor capital stock. We find a steady-state value of $k$, $\overline{k}$, and then use this to compute the balanced-growth rate of asset returns $\overline{R}=(1-\alpha)\overline{k}^{-\alpha}$.
@@
Collecting all of these results gives us a relative consumption ratio that depends on whether or not inculcation and gift giving will occur in equilibrium, which can be obtained by using equilibrium values of $\overline{R}$ and $\eta$ in Equation (46) in Proposition 7:
@@
Increases in the rate of population growth increase the relative well-being of the elderly, whether or not there are gifts involved. In the high-capital intensity world, in which there is no gift-giving and inculcation, it is because population growth increases the relative scarcity of capital, which increases rates of return on accumulated assets, benefiting the elderly. This is also true in the second case, but the effect is augmented by another aspect of how gifts interact with population growth in the model - when there are more young people relative to elderly, there are more gifts to go around.
The relative intensity of capital in production, $\frac{1-\alpha}{\alpha}$, plays a key role in influencing the relative treatment of the elderly. The underlying reasons are similar to those described in the context of the model specification developed in Section 4.2. When production is more labor-intensive, the relative treatment of the elderly tends to decline, but at some point, inculcation and acculturation of elderly treatment come into play, mitigating this trend. Put another way, when income depends mainly on labor and not accumulated assets, there is an incentive to develop cultural practices facilitating transmission of income to the elderly, as there are no other ways to accumulate. The flip side is that when production is more asset-intensive, the relative treatment of the elderly may improve while simultaneously acculturation to treat the elderly well wanes. The full ramifications of the elderly treatment model are shown in Figure 3, which shows the relative consumption of the elderly $\frac{c_{e}}{c_{m}}$ in (56) as a function of the capital intensity of production. One can see from the figure the same sort of emergent nonlinear, U-shaped arrangement between relative elderly treatment as occurred before in the property-rights-based model of Section 4.2.
@@ 4.5.2 Full model solution
The results of the previous Section applied when the sole source of income for the elderly is previously accumulated assets. When the elderly can save, but also have other sources of income deriving from their labor available, the resulting model is more complex because not only does elderly labor income provide sustenance when elderly, but it also raises lifetime income, which influences incentives both to give gifts and to preserve wealth for future consumption. These effects can make it difficult to accurately state how changes in model parameters influence inculcation, accumulation, and relative treatment of the elderly. In spite of this, one can still solve the model assuming that all model variables are along a balanced growth path. With $\tau_{e},\gamma_{e}>0$, the steady-state relationships corresponding to (42) and (43) in Proposition 6 can be written as:
@@
From (57), we see that as $\tau_{e}$ increases, meaning that the elderly have a relatively larger source of income, $R$ increases in an equilibrium with a positive amount of inculcation ($\eta>0$). This is because when a future source of income is available, the incentives to save are deadened. For the same reason, increases in $\tau_{e}$ lessen inculcative investments, as is evident from (58).
@@
Using (60) and $\eta=0$ in (46), we find that relative consumption becomes.
@@ 5 Discussion and Conclusions
The model presented in this paper serves to reinforce previous work on the relationship between intergenerational transfers and their role in delivering support to the elderly. But our model also allows us to discuss more thoroughly the features of the extensive cross-cultural research that has been done on the treatment of the elderly. Our analysis spans not only relatively recent phenomena such as the demographic transition, but also helps explain earlier changes in the treatment of the elderly across the entire spectrum of levels of development. Our model extends to matters such as the development of property rights and how property rights interact with culture and elderly treatment. Our model focuses attention squarely on elderly treatment, not on intermediate features of the economy such as the occurrence and direction of transfers.
@@
Our model demonstrates that the relationship between things such as ownership, accumulation, and production and the treatment of the elderly is intrinsically complex, even as we are able to lend theoretical support for features of the ethnographic record such as the ‘curvilinear hypothesis.’ Moreover, with our model we are also able to discuss how the importance of accumulation in production impacts the elderly. Our findings here are also subtle; the possibility of accumulation may initially have an adverse impact on the relative status of the elderly. However, increasing the importance of accumulation may eventually function to the relative advantage of the elderly. Indeed, our results lend further support to the hypothesis that there is a nonlinear relationship between the complexity of familial structure and development (Nimkoff and Middleton, 1960; Blumberg and Winch, 1972). This observation in fact echoes the general trend observed by (Murdock and Provost, 1973) - that the most culturally sophisticated societies tend to lie at the middle spectrum of development, who go so far to say that “This essentially bimodal or curvilinear distribution is inconsistent with any unilinear interpretation of social development,” (Murdock and Provost, 1973, p. 392) our model suggests why such a pattern might exist, at least in regards elderly treatment and coincident transmission of culture. [^22] This tendency has been described in the context of land inheritance rules by Baker and Miceli (2005) and in marriage institutions (Baker and Jacobsen, 2007a, b).
There are several possible avenues for further research. Some are suggested by our theory, while others derive from aspects of the ethnographic record that we did not treat in detail for reasons of brevity. One such topic is the relationship between family structure, technical change, and the treatment of the elderly. To the extent that gift giving can imply a closer family, our model informs some of these issues, but they remain an interesting avenue for future research, along with the relationship between elderly treatment, marriage, and fertility. [^23] These issues are partially explored in Cigno et al. (2020) and in Bisin and Verdier (2000). We also have not addressed gender differences in elderly treatment, and to be sure, documenting such differences is a point of emphasis of Simmons (1970). There are also potential complexities of the problem of inculcation and cultural transmission more generally that could be further explored. To the extent that inculcation and elderly treatment are behavioral norms, there may be important dynamic and public-goods aspects to decisions that govern these phenomena, some of the complexities of which figure prominently in Bowles (1998). Our model has also focused only on inter-societal relative well-being of the elderly, but a remaining issue to be explored is inequality among the elderly, where poverty differences and inequality among the elderly may be related to cultural transmission and gift giving as well. Our hope is that our general approach will be fruitful when applied to these additional topics.
Appendix A Proof of Proposition 4

### v2

@@ 4.4 Savings, Capital Accumulation, and Inculcation
We now analyze a version of the model where assets are accumulated through conscious savings decisions, exploring how this interacts with inculcation and gifts to the elderly. This approach also connects inculcative investment with savings and aligns the exposition with classical growth models featuring technological progress.
@@
The solution of (35) presents two challenges. First, a solution must take into account that current middle-aged agents’ future gifts depend on current inculcative investment through the parameter $\eta_{t+1}$, and in an equilibrium solution to the model, beliefs about this dependence must be correct. Second, a solution must admit the possibility of solutions in which $\eta^{*}_{t}=0$. The subgame-perfect Nash equilibrium solution to the problem (35) has a structure in which agents give gifts in proportion to the present discounted value of lifetime income, with this proportion depending on $\eta_{t}$. For the time being, we have bypassed the issue of the determination of $\eta_{t}^{*}$ and skip to the form of the gift-giving function, which is given in Proposition 4:
@@
See Appendix C. ∎
Equation (37) in Proposition 5 describes the total amount middle-aged agents wish to allocate to future consumption. Expenditures on savings and inculcation add up to the fraction of income a middle-aged agent wishes to save, which depends upon current income less discounted future income. Equation (38) in Proposition 5 requires that the rates of return on the different means of transferring current wealth to the future in (37) be equilibrated. The rate of return on inculcative investments and conventional savings must be equal; otherwise, middle-aged agents would not employ both savings instruments. As we maintain the assumption that $d_{t}$ is proportional to income ($d_{t}=\delta y_{mt}$), we can interpret (38) to mean that returns to assets such as capital must be equal in equilibrium to the unit return on inculcative investment.
The question remains whether an equilibrium with gifts and inculcation is sustainable; i.e., whether $\eta_{t}>0$ in equilibrium. We need to round out the model to see how $\eta_{t}$, $s_{t}$, and $R_{t}$ depend on the underlying nature of the economy, as they are quantities determined in equilibrium and require specification of the production side of the economy.
@@ 4.4.1 Production and Economic Growth
We again use a Cobb-Douglas production technology in which production occurs according to the aggregate production function $Y_{t}=L_{t}^{\alpha}K_{t}^{1-\alpha}$, where $K_{t}$ now denotes the total capital stock and as before $L_{t}$ denotes total effective labor supply:
@@
We assume that $A_{it},i=e,m$ grows at a constant exogenous rate $a$, while the population grows at a rate $n$. Constant population growth means that $N_{mt}=(1+n)N_{et}$, and that $\tfrac{1}{1+n}$ is a measure of the relative size of the elderly population.
We use (39) and the production technology to obtain the following expressions for middle-aged labor income, elderly labor income, and the gross rate of return on capital, assuming that all resources are paid their marginal products, where $k_{t}=\frac{K_{t}}{L_{t}}$ is capital per effective unit of labor:
@@
We also define some new terms:
@@
See Appendix E. ∎
Equation (43) describes the dynamics of inculcation but also determines whether there will be any inculcative investment at all in equilibrium. When the right-hand side of (43) is nonpositive, $\eta_{t+1}=0$ in equilibrium. This possibility depends on the rates of population growth and technological progress; in particular, a higher rate of technological progress makes $\eta=0$ more likely. Higher growth rates increase the demand and hence returns on capital, making inculcation less desirable. Also making $\eta=0$ more likely in equilibrium are higher inculcative costs, as measured by $\delta$, and greater direct participation in the economy by the elderly (larger values for both $\tau_{e}$ and $\gamma_{e}$).
A value of the relative consumption ratio of the elderly $\frac{c_{et}}{c_{mt}}$ can be found using the equilibrium values of savings, inculcation, and gifts. In Appendix E, we describe the calculations that lead to Proposition 7:
@@
Proposition 7 indicates that the relative consumption ratio depends on relative lifetime incomes, and therefore the overall growth rate of the economy. The proposition also suggests that equilibrium asset returns, $R$, and the magnitude of $\eta$ are important in driving relative consumption. This is particularly true along a balanced growth path, as in a balanced growth equilibrium, incomes grow at a rate $a$, and asset returns are constant. Along a balanced growth path, (44) reduces to:
@@
Equation (46) therefore indicates that parameters such as $\alpha$, the labor intensity of production, $n$, the growth rate of population, and $\delta$, the per-unit-income costs of inculcation will influence the relative consumption of the elderly by exerting an impact on the level of inculcation $\eta$ and equilibrium asset returns $R$. Of course, $\eta$ and $R$ are somewhat complicated functions of model parameters, as discussed in Section 4.5.2.
@@ 4.5.1 Asset income as the sole source of elderly income
Equation (51) shows how the possibility of inculcation and gifts impacts the canonical overlapping-generations model with population and technological progress, as developed in Acemoglu (2009). But for the $\delta\eta_{t+1}$ term in (51), the accumulation condition is a standard capital accumulation condition in growth models, so the possibility of inculcative investment reduces the amount of capital per effective unit of labor, which will increase equilibrium asset return rates. Dividing (50) by (51) and solving the result for $\eta_{t+1}$ gives a time-independent solution for the inculcative parameter $\eta$:
@@
Equation (52) indicates that a larger $\delta$ reduces the equilibrium size of $\eta^{*}$. An increase in the relative intensity of capital in production, measured by $\frac{1-\alpha}{\alpha}$, also negatively impacts $\eta$. Societies that employ more capital-intensive means of production, other things equal, will have lower degrees of inculcation.
If parameters are configured so that $\eta^{*}\leq 0$, in equilibrium $\eta=0$, and there is no inculcative investment. The elderly ensure their livelihood as they would in the typical overlapping generations model: through prior accumulation of productive assets. Setting (52) to zero and solving for $\alpha$ gives an interpretation of (52) as a condition for inculcation to occur in terms of the relative capital intensity of production:
@@
This threshold depends positively on the costs of inculcation, and on $\beta$, which influences the relative weight placed on consumption in old age in utility. If (53) is not satisfied and no inculcation occurs, we recover the law of motion for the capital stock by using $\eta=0$ in (51). In forming the relative ratio of consumption, we find it most straightforward to solve the model in terms of $R$, equilibrium asset returns, rather than in terms of $k$, the equilibrium per unit of labor capital stock. We find a steady-state value of $k$, $\overline{k}$, and then use this to compute the balanced-growth rate of asset returns $\overline{R}=(1-\alpha)\overline{k}^{-\alpha}$.
@@
Collecting all of these results gives us a relative consumption ratio that depends on whether or not inculcation and gift giving will occur in equilibrium, which can be obtained by using equilibrium values of $\overline{R}$ and $\eta$ in Equation (46):
@@
Population growth increases the relative well-being of the elderly, whether or not gifts are involved. In a high-capital-intensity setting without gift-giving or inculcation, this occurs because population growth raises the relative scarcity of capital, boosting returns on accumulated assets and benefiting the elderly. In cases with gift-giving, this effect is amplified by the greater number of young people, leading to more gifts being distributed to the elderly.
The capital intensity of production, $\frac{1-\alpha}{\alpha}$, significantly influences the relative treatment of the elderly. As in the model from Section 4.2, more labor-intensive production tends to lower the relative treatment of the elderly. However, cultural practices such as inculcation and acculturation emerge to mitigate this decline when income primarily depends on labor rather than accumulated assets, providing a way to transfer income to the elderly. Conversely, in more asset-intensive production, the elderly’s relative treatment improves due to asset income, but cultural acculturation to support the elderly diminishes.
Figure 3 illustrates this dynamic, showing the relative consumption of the elderly, $\frac{c_{e}}{c_{m}}$, from (56) as a function of capital intensity. The figure reveals a nonlinear, U-shaped relationship between elderly treatment and capital intensity, similar to the pattern seen in the property-rights model of Section 4.2.
@@ 4.5.2 Full model solution
When the elderly can save and also have a sources of labor income, the model is more complex because elderly labor income provides sustenance and also raises lifetime income. This makes it difficult to accurately state how changes in model parameters influence inculcation, accumulation, and relative treatment of the elderly. With $\tau_{e},\gamma_{e}>0$, the steady-state relationships corresponding to (42) and (43) in Proposition 6 can be written as:
@@
From (57), we see that as $\tau_{e}$ increases, meaning that the elderly have a relatively larger source of income, $R$ increases in an equilibrium with a positive amount of inculcation ($\eta>0$). When a future source of income is available, the incentives to save are deadened. Similarly, increases in $\tau_{e}$ lessen inculcative investments, as is evident from (58).
@@
Using (60) and $\eta=0$ in (46), we find that relative consumption becomes:
@@ 5 Discussion and Conclusions
The model in this paper builds on prior research into intergenerational transfers and their role in supporting the elderly while allowing for a deeper exploration of cross-cultural findings on elderly treatment. Our analysis spans both recent phenomena, such as the demographic transition, and historical shifts in elderly treatment across varying levels of technological sophistication. By incorporating the interplay between property rights, culture, and elderly treatment, the model centers attention on elderly well-being rather than intermediate economic features like transfer patterns.
@@
The model highlights the complexity of relationships between ownership, accumulation, production, and elderly treatment, providing theoretical support for observed ethnographic patterns including the ‘curvilinear hypothesis.’ It also addresses how accumulation in production impacts the elderly, showing that while initial accumulation may disadvantage them, increasing its importance can eventually benefit their relative status. These findings align with the hypothesis of a nonlinear relationship between familial complexity and degree of modernization (Nimkoff and Middleton, 1960; Blumberg and Winch, 1972). They echo Murdock and Provost (1973), who observed that the most culturally advanced societies often fall at the mid-spectrum of technological complexity, noting, “This essentially bimodal or curvilinear distribution is inconsistent with any unilinear interpretation of social development” (Murdock and Provost, 1973, p. 392). Our model offers insights into why such a pattern might arise, particularly regarding elderly treatment and cultural transmission. [^13] Similar patterns are discussed in the context of land inheritance by Baker and Miceli (2005) and marriage institutions by Baker and Jacobsen (2007a, b).
Several avenues for future research emerge from this work. One is the interplay between family structure, technological change, and elderly treatment. While our model touches on these issues through its focus on gift-giving and familial ties, deeper exploration of the connections between elderly treatment, marriage, and fertility is warranted. [^14] See related discussions in Cigno et al. (2020) and Bisin and Verdier (2000). Gender differences in elderly treatment also deserve attention, as emphasized by Simmons (1970). Furthermore, the dynamics of inculcation and cultural transmission, including their public-goods and behavioral aspects, remain ripe for investigation (Bowles, 1998). Finally, while our focus has been on inter-societal well-being of the elderly, future research could address inequality among the elderly, examining how poverty and disparities relate to cultural transmission and gift-giving practices. We hope this framework will inspire further inquiry into these important topics.
Appendix A Data Appendix
We describe the data underlying the analysis presented in Tables 1 and 2 of the text and also provide an overview of the data in Simmons (1970). [^15] All data are derived directly from Simmons (1970), and are available on request from the authors. Simmons developed variables to measure the presence and importance of material, social, and cultural characteristics across 71 cultures, including hunters, pastoralists, nomads, and agriculturalists. Features were coded as $+$ (very important), $=$ (somewhat important), -- (present but unimportant), or 00 (absent). We recoded nonmissing values numerically, ranging from 3 (very important) to 0 (absent).
Simmons’s data includes 224 features, with some values specific to women, men, or the elderly. However, missing values and vague variable definitions make formal analysis beyond Simmons’s cross-tabulations challenging and potentially misleading. Therefore, we adopted a descriptive approach, compiling summary indices from his data. The evidence presented in the text is based on these indices, whose construction we now discuss.
Table A11 describes the variables that we judge as capturing the market-openness of the economic environment present; specifically, the degree to which property rights, exchange, and market orientation are features of a society.
Mean SD Min Max N
Trade presence 2.33 0.75 1 3 67
Money in use 1.70 1.07 0 3 57
Private ownership of objects 2.60 0.71 1 3 63
Private ownership of land 1.37 1.19 0 3 54
Male property rights 2.58 0.72 1 3 53
Female property rights 1.56 0.98 0 3 32
Communal ownership of land 2.21 0.76 0 3 56
Table A1: Variables included in the openness of economy index, which captures the degree to which a market-oriented economy might be present in each society
Our market openness index is constructed by simply adding and subtracting characteristics. While more complex methods like principal components could be used, experimentation showed little benefit compared to this transparent approach. We sum the first six characteristics in Table A11 and subtract the last (communal land ownership). Summary statistics for the index are in Table A88.
Simmons also notes how actively the elderly participate in aspects of the economy. The variables summarized in Table A22 include jobs that the elderly do in society that we judge to directly support the production of physical output in a society, or perform functions that free other members of society to engage in direct production.
Mean SD Min Max N
Male traders 2.35 0.75 1 3 20
Female traders 1.33 0.52 1 2 6
Male provision of infant services 1.79 0.93 0 3 24
Female provision of infant services 2.49 0.83 0 3 47
Male education of young 2.83 0.45 1 3 40
Female education of young 2.86 0.35 2 3 36
Male nurses 2.90 0.31 2 3 20
Female nurses 2.67 0.72 1 3 15
Males aid with agriculture 1.82 1.38 0 3 17
Females aid with agriculture 2.15 1.31 0 3 20
Males aid with hunting 2.53 0.62 1 3 17
Females aid with hunting 2.67 0.52 2 3 6
Males aid with herding 1.29 1.44 0 3 14
Females aid with herding 0.83 1.27 0 3 12
Males aid with collecting 2.50 0.71 2 3 2
Females aid with collecting 2.67 0.52 2 3 6
Males aid in household 2.55 0.52 2 3 11
Females aid in household 2.84 0.37 2 3 32
Males make toys and tools 2.36 0.63 1 3 14
Females make toys and tools 2.17 1.17 0 3 6
Male elderly fees for services 2.80 0.41 2 3 25
Female elderly fees for services 2.56 0.62 1 3 18
Table A2: Variables included in market development and sophistication index
From these variables, we form what we call a direct production index by simply adding all the variables in the table, treating missing values as zeros. The resulting index is summarized in Table A88 and is designed to capture something about the degree to which the elderly contribute to material life in each culture.
The next set of variables we use to construct an index of the degree to which the elderly perform important ritualistic or social functions. These variables are summarized on table A33; the sum of the variables comprises our index of social functions, also summarized in table A88.
Mean SD Min Max N
Male rites of initiation 2.40 1.00 0 3 25
Female rites of initiation 1.90 0.94 0 3 21
Males perform marriage rites 2.77 0.43 2 3 31
Females perform marriage rites 2.43 0.60 1 3 21
Males perform funereal rites 2.60 0.76 0 3 25
Females perform funereal rites 2.54 0.51 2 3 26
Male priesthood 2.86 0.36 2 3 35
Female priesthood 1.85 0.69 0 3 13
Male shamans 2.86 0.39 1 3 59
Female shamans 2.49 0.69 0 3 47
Male decorators 2.33 0.58 2 3 3
Female decorators 2.67 0.50 2 3 9
Table A3: Social functions performed by the elderly
The variables summarized in Table A4 4 describe the degree to which the elderly provide advice, store important social information, or perform other functions that require knowledge. We computed an index of knowledge provision based on just summing up these variables.
Mean SD Min Max N
Males provide expert advice 2.84 0.37 2 3 25
Females provide expert advice 2.71 0.49 2 3 7
Male stores of social information 2.98 0.14 2 3 54
Female stores of social information 2.59 0.50 2 3 27
Males lead festivals 2.69 0.52 1 3 42
Females lead festivals 2.50 0.64 1 3 28
Table A4: Knowledge provision by the elderly
The variables in Table A55 describe the degree to which the elderly hold important titles in society, function as elders, judges, or chiefs. Our index of titular duties is created as a simple sum of these variables.
Mean SD Min Max N
Male chiefs 2.37 0.71 1 3 62
Female chiefs 0.35 0.61 0 2 31
Male council of elders 2.60 0.65 0 3 57
Male judges 2.56 0.69 0 3 45
Female judges 0.62 0.87 0 2 13
Table A5: Titular functions of the elderly
Various variables catalog the degree to which the elderly are respected, supported by the community, or more generally viewed by society. These variables are collected in Table A66, and we collect these variables into an index that is a simple sum of these variables.
Mean SD Min Max N
Male elderly respected 2.81 0.47 1 3 63
Female elderly respected 2.29 0.81 0 3 51
Male elderly community support 2.09 0.84 0 3 33
Female elderly community support 1.97 0.81 0 3 30
Male elderly family support 2.80 0.48 1 3 60
Female elderly family support 2.58 0.57 1 3 57
Male son-in-law elderly support 1.82 1.05 0 3 22
Female son-in-law elderly support 2.06 1.06 0 3 16
Male elderly seen as distinguished 2.71 0.61 1 3 14
Female elderly seen as distinguished 3.00 . 3 3 1
Table A6: Positive treatment practices towards the elderly
Some variables actually measure the degree to which the elderly are pictured as having magical powers and related beliefs. We gather all of these variables into an index, which is a simple sum of the variables in Table A67.
Mean SD Min Max N
Males glorified as daimons 2.64 0.54 1 3 36
Females glorified as daimons 2.44 0.65 1 3 25
Males glorified as heros 2.77 0.43 2 3 39
Females glorified as heros 2.55 0.68 1 3 31
Males friends of children 2.62 0.52 2 3 8
Females friends of children 2.45 0.82 1 3 11
Table A7: Inculcated beliefs about the elderly
Table A88 contains a summary of all indices that we have created. These indices are used to calculate the correlations in Table 1 in the text.
Mean SD Min Max N
Openness of economy index 6.49 4.27 -3 14 71
Elderly contribution to direct production index 6.99 4.35 0 17 71
Contribution to social production index 8.72 3.71 0 18 71
Knowledge and advice provision index 3.55 2.14 0 8 71
Titular duties of elderly index 3.02 1.44 0 6 71
Positive treatment of elderly index 5.99 2.68 0 13 71
Positive attitudes to elderly inculcation index 2.75 2.22 0 9 71
Elderly social functions index 5.70 2.98 0 14 71
Total elderly contributions in production index 19.25 8.88 2 40 71
Table A8: Summary statistics for constructed indices used in Table 1.
Appendix B Proof of Proposition 4